OpenStudy (anonymous):

find equation of line through points (4, -5) and (1, -1) PLEASE HELP FINAL EXAM TOMORROW

OpenStudy (aaliyahfairgood):

first, find the slope m=y2-y1/x2-x1 m=(-1-(-5))/(1-4) m=4/-3 then use slope int.form y-y1=m(x-x1) y-(-5)=-4/3(x-4) y+5=-4/3x+16/3 -5(1/3) on both sides Multiply by 1/3 to get a common denominator for 16/3 y=-4/3x+1/3

OpenStudy (anonymous):

sorry i was trying to work it out. I'm stuck at y+5=-4/3+16/3

OpenStudy (aaliyahfairgood):

well you multiply negative 4 thirds by negative four and get 16/3 then you have to subtract 5 to get y alone. To do that you have to find a common denominator so you can subtract 5 into 16/3

OpenStudy (aaliyahfairgood):

so look at 5 as a fraction 5/1 so (16/3)-(5/1)

OpenStudy (aaliyahfairgood):

multiply 5 by 1/3 and you have 16/3 minus 12/3 wow I messed up but the result is 4/3 so actually, y=-4/3x+1/3is wrong the final answer is y=-4/3x+4/3 sorry about that